OpenStudy (anonymous):

The larger of two numbers is 5 less than twice the small number. the sum of the two numbers is 31. Find the two numbers

OpenStudy (anonymous):

So, say larger number is L, smaller is S. Then: L = 2S - 5 And: L + S = 31 ----> L = 31 - S 31-S = 2S -5 36 = 3S S = 12 Then, L = 2(12) - 5 = 24-5 = 19. So numbers are 19 and 12.